OpenStudy (anonymous):

what does the geometric sequance formula look like? and can the geometric (series) formula find the sum of a geo sequance problem??

OpenStudy (anonymous):

where is the formula?

OpenStudy (amistre64):

there are 2 versions of it that i can recall \[A_{n+1}=A_n*r\]\[A_n=A_1*r^{n-1}\]

OpenStudy (anonymous):

oh, okay. can these formulas solve lets say for example: What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms?

OpenStudy (amistre64):

not as is; they just allow us to find the terms of the sequence itself a series (adding up the terms of the seq) goes by another formula \[a\frac{1-r^n}{1-r}\]

OpenStudy (amistre64):

a, in this case, is the first term

OpenStudy (anonymous):

so that would be the formula I would use for that problem?

OpenStudy (amistre64):

if you want to add up all the terms; then yes

OpenStudy (amistre64):

a = 8, r=-16/8, and n=15 right?

OpenStudy (anonymous):

yep, and then I just pulg them in for the terms.

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

would this be my answer? 87,384?

OpenStudy (anonymous):

lol thank you, at least yu took the time to help me :)

OpenStudy (amistre64):

youre welcome, and good luck :)